OpenStudy (anonymous):

sum of the infinite geometric series (1/2)+(1/4)+(1/8)+(1/16)+...

sam (.sam.):

\[a_1+a_2+a_3+a_4+...\] \[r=\frac{n_2}{n_1}\] \[S_\infty=\frac{a_1}{1-r}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

how do you find r though?

sam (.sam.):

Use 1/4 divide 1/2 gets you r

OpenStudy (anonymous):

picture above shows the answer without any computation just thought i would mention it

sam (.sam.):

Lol how do you use that diagram to answer

OpenStudy (anonymous):

fill it in you can see that it will fill up the whole unit square, the further you go so the answer is one

OpenStudy (anonymous):

also you might note that in binary \[\frac{1}{2}+\frac{1}{4}+\frac{1}{8}+...=0.11111...\]

OpenStudy (anonymous):

ohhhh i see thank you both :D

OpenStudy (anonymous):

and just like if you have base ten \(0.999...=1\) so in base two \(0.11111...=1\)

sam (.sam.):

ah

OpenStudy (anonymous):

but really you are supposed to do what @.Sam. said \[\frac{\frac{1}{2}}{1-\frac{1}{2}}\]

OpenStudy (anonymous):

you might also note that in base 3 \[\frac{1}{3}+\frac{1}{3^2}+\frac{1}{3^3}+...=0.1111111\]

OpenStudy (anonymous):

that's interesting :D thank you!!

OpenStudy (anonymous):

yw

sam (.sam.):

I like your different approach